### v1

# Feynman Integral Reduction without Integration-By-Parts

Abstract
We present an interesting study of Feynman integral reduction that does not employ integration-by-parts identities. Our approach proceeds by studying the equivalence relations of integral contours in the Feynman parameterization. We find that the integration contour can take a more general form than that given by the Cheng-Wu theorem. We apply this idea to one-loop integrals, and derive universal reduction formulas that can be used to efficiently reduce any one-loop integral. We expect that this approach can be useful in the reduction of multi-loop integrals as well.
1 Introduction
Scattering amplitudes and Feynman integrals are core components of perturbative quantum field theories. Developing efficient computational methods is crucial for advancing cutting-edge phenomenological applications. To compute multi-loop scattering amplitudes, it is often necessary to handle a large number of complicated Feynman integrals. By exploiting the linear relations among these integrals, they can be expressed as linear combinations of a finite set of master integrals (MIs). This process is known as integral reduction, which significantly reduces the computational complexity. Integral reduction is also a key step in the method of differential equations Kotikov:1990kg ; Kotikov:1991pm ; Remiddi:1997ny ; Gehrmann:1999as for evaluating the MIs.
Currently, the standard method for integral reduction is the Laporta algorithm Laporta:2000dc for solving the integration-by-parts (IBP) identities Tkachov:1981wb ; Chetyrkin:1981qh of Feynman integrals. Many program packages implementing this algorithm are available, including FIRE Smirnov:2023yhb , LiteRed Lee:2013mka , Reduze vonManteuffel:2012np and Kira Klappert:2020nbg . Traditionally, the IBP method has been developed mainly in the momentum representation. They can be formulated in other representations as well, such as in the Baikov representation Lee:2014tja ; Larsen:2015ped ; Bendle:2019csk ; Chen:2022jux or in the Feynman parameterization Lee:2014tja ; Chen:2019mqc ; Chen:2019fzm ; Chen:2020wsh ; Chen:2024xwt . For complicated multi-loop integrals, the system of IBP equations can become very large, and it is rather time-consuming or even impractical to solve them. Recently, it has been proposed to generate a smaller set of IBP relations by using the algebraic geometry based method in NeatIBP Wu:2023upw , or by searching for a block-triangular system in Blade Guan:2024byi . In addition to the IBP method, there exist other reduction techniques as well, e.g., Passarino-Veltman (PV) tensor reduction Passarino:1978jh and its improvements using auxiliary vectors Feng:2021enk ; Hu:2021nia ; Feng:2022rwj ; Feng:2022uqp ; Feng:2022iuc ; Feng:2022rfz ; Li:2022cbx , Ossola-Papadopoulos-Pittau (OPP) method Ossola:2006us ; Ossola:2007bb ; Ellis:2007br , unitarity cut method Bern:1994zx ; Bern:1994cg ; Bern:1997sc ; Britto:2004nc ; Britto:2005ha ; Britto:2006sj ; Anastasiou:2006jv ; Anastasiou:2006gt ; Britto:2006fc ; Britto:2007tt ; Britto:2010um , generating functions Feng:2022hyg ; Guan:2023avw ; Hu:2023mgc ; Li:2024rvo , et al.. We will not go into details of these methods.
An alternative way to formulate the IBP relations is the so-called intersection theory Mizera:2017rqa ; Mastrolia:2018uzb ; Frellesvig:2019kgj ; Frellesvig:2019uqt ; Mizera:2019vvs ; Mizera:2020wdt ; Frellesvig:2020qot ; Caron-Huot:2021xqj ; Caron-Huot:2021iev ; Chestnov:2022alh ; Fontana:2023amt ; Brunello:2023rpq , where a Feynman integral is regarded as a pairing between a differential form and an integration contour. The IBP relations are formulated as the equivalence relations among differential forms, which live in a so-called twisted cohomology group. The integral reduction can then be performed by calculating the intersection numbers between a pair of differential forms. Such an approach has been extensively developed in the Baikov representation Baikov:1996iu ; Lee:2010wea of Feynman integrals, and recently has been developed for the Feynman parametrization as well Lu:2024dsb .
Within the framework of intersection theory, the equivalence relations can also be established among integration contours. The equivalence classes of contours form a twisted homology group. In principle, these equivalence relations can also be employed for integral reduction, but this approach has not been developed so far.
In this work, we initiate a study that exploits the equivalence of integration contours for integral reduction, based on the Feynman parameterization. One outcome of our study is an improvement of the Cheng-Wu theorem Cheng:1987ga , such that the delta-function in the Feynman parameterization can be modified to a more general form. This essentially corresponds to modifying the integration contour. We apply this to one-loop integrals and find that, by splitting the contour and further transforming each part of the contour, we can identify each part with a Feynman integral that is simpler than the original one. As a result, we can construct recursive reduction formulas purely by dealing with integration contours, without solving IBP relations. This approach does not generate any redundant information hidden in the IBP relations, and is therefore highly efficient.
The paper is organized as follows. In Section 2, we introduce the equivalence relations of integral contours in Feynman parameterization and use several simple examples to demonstrate our reduction method. In Section 3, we present our general method for one-loop integral reduction, and provide recursive formulas that can be easily implemented in computer algebra. In Section 4, we demonstrate our method through several examples. In Section 5, we provide a summary and explore the potential extension of this method to higher loops, discussing the new challenges that may arise in future applications.
2 Domain of integration and reduction of Feynman integrals
2.1 Feynman parametrization and equivalence classes of integration domains
An $L$-loop Feynman integral is defined by
$I(\bm{\nu}_{n})=e^{\epsilon\gamma_{E}L}\int\frac{\mathrm{d}^{d}k_{1}}{i\pi^{d/2}}\cdots\frac{\mathrm{d}^{d}k_{L}}{i\pi^{d/2}}\,\frac{1}{D_{1}^{\nu_{1}}\cdots D_{n}^{\nu_{n}}}\,,$ (1)
where $\bm{\nu}_{n}\equiv\{\nu_{1},\cdots,\nu_{n}\}$, and $D_{i}$ are propagator denominators or irreducible scalar products. One can convert the above momentum representation into integrals over Feynman parameters. There are many variants of the Feynman parametrization, and one of them was introduced in Chen:2019mqc . It takes the form (assuming all $\nu_{j}>0$):
$I(\bm{\nu}_{n})=C({\bm{\nu}}_{n})\int_{0}^{\infty}\left(\prod_{j=0}^{n}\alpha_{j}^{\nu_{j}-1}\mathrm{d}\alpha_{j}\right)\left(\alpha_{0}\,\mathcal{U}+\mathcal{F}\right)^{\lambda_{0}}\,\delta\left(1-\sum_{j\in\mathcal{S}}\alpha_{j}\right)\,,$ (2)
where $\mathcal{S}$ is a non-empty subset of $\{0,1,2,\cdots,n\}$, $\lambda_{0}\equiv-d/2$, $\nu_{0}\equiv-\nu-(L+1)\lambda_{0}$, $\nu\equiv\sum_{j=1}^{n}\nu_{j}$, and the prefactor is given by
$C({\bm{\nu}}_{n})\equiv{(-1)}^{\nu}e^{\epsilon\gamma_{E}L}\frac{\Gamma(-\lambda_{0})}{\prod_{i=0}^{n}\Gamma(\nu_{i})}\,.$ (3)
$\mathcal{U}$ and $\mathcal{F}$ are the so-called Symanzik polynomials. We denote
$\alpha_{1}D_{1}+\cdots+\alpha_{n}D_{n}\equiv\sum_{i,j=1}^{L}M_{ij}\,k_{i}\cdot k_{j}-2\sum_{i=1}^{L}k_{i}\cdot Q_{i}-J+i0\,,$ (4)
where $Q_{i}$ are combinations of external momenta. The two Symanzik polynomials can then be written as
$\mathcal{U}=\det(M)\,,\quad\mathcal{F}=\det(M)\left[\sum_{i,j=1}^{L}M^{-1}_{ij}\,Q_{i}\cdot Q_{j}-J-i0\right]\,.$ (5)
From the above expressions, it is clear that $\mathcal{U}$ and $\mathcal{F}$ are homogeneous polynomials of degree $L$ and $L+1$ in the variables $\{\alpha_{1},\cdots,\alpha_{n}\}$, respectively.
Note that the representation (2) can be applied to the cases where some $\nu_{j}\leq 0$ as well. For that we can introduce a regularization $\alpha_{j}^{\rho}$ into the integration measure. After performing the reduction, one takes the limit $\rho\to 0$ in the end. In the following, we will assume that such regulators are implicitly applied when necessary.
The fact that one can freely choose a subset of Feynman parameters appearing in the $\delta$-function in Eq. (2) follows from the so-called Cheng-Wu theorem Cheng:1987ga . We now note that Eq. (2) can actually be recasted into a more general form:
$I(\bm{\nu}_{n})=C({\bm{\nu}}_{n})\int_{0}^{\infty}\left(\prod_{j=0}^{n}\alpha_{j}^{\nu_{j}-1}\mathrm{d}\alpha_{j}\right)\left(\alpha_{0}\,\mathcal{U}+\mathcal{F}\right)^{\lambda_{0}}\,\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\,,$ (6)
where $\mathcal{X}$ and $\mathcal{Y}$ are non-negative homogeneous functions of degree 00 and $1$ in the variables $\{\alpha_{j}\}$, and we also require that the equation $\mathcal{X}-\mathcal{Y}=0$ has only one solution for any variable $\alpha_{j}$ in the whole integration domain. The original representation (2) then corresponds to the specific choice $\mathcal{X}=1$ and $\mathcal{Y}=\sum_{j\in\mathcal{S}}\alpha_{j}$. Note that by choosing $\mathcal{X}=1$ and $\mathcal{Y}=\alpha_{0}$, we can integrate out $\alpha_{0}$ to arrive at the so-called Lee-Pomeransky (LP) representation Lee:2013hzt :
$I(\bm{\nu}_{n})=C({\bm{\nu}}_{n})\int_{0}^{\infty}\left(\prod_{j=1}^{n}\alpha_{j}^{\nu_{j}-1}\mathrm{d}\alpha_{j}\right)\left(\mathcal{U}+\mathcal{F}\right)^{\lambda_{0}}\,.$ (7)
We now show why Eq. (6) is equivalent to Eq. (2). The integral (6) can be regarded as the integration of a differential $n$-form on the $n$-dimensional hypersurface determined by $\mathcal{X}-\mathcal{Y}=0$:
$S_{n}\equiv\left\{\left(\alpha_{0},\cdots,\alpha_{n}\right)\in\mathbb{R}^{n+1}\big{|}\mathcal{X}-\mathcal{Y}=0\,\&\,\alpha_{j}\geq 0,\forall j\right\}.$ (8)
The differential $n$-form can be written as $f\,\omega$, where
$\omega\equiv\sum_{j=0}^{n}(-1)^{j}\,\alpha_{j}\,\mathrm{d}\alpha_{0}\wedge\ldots\wedge\widehat{\mathrm{d}\alpha_{j}}\wedge\ldots\wedge\mathrm{d}\alpha_{n}\,,$ (9)
where the hat indicates that the corresponding factor is omitted, and
$f\equiv C(\bm{\nu}_{n})\left(\prod_{j=0}^{n}\alpha_{j}^{\nu_{j}-1}\right)\left(\alpha_{0}\,\mathcal{U}+\mathcal{F}\right)^{\lambda_{0}}\,.$ (10)
With the above definitions, the integral in Eq. (6) can be written as
$I(\bm{\nu}_{n})=\int_{S_{n}}f\,\omega\,.$ (11)
To show that the above relation is true, we use the fact that when restricted on $S_{n}$,
$\sum_{j=0}^{n}\mathrm{d}\alpha_{j}\frac{\partial(\mathcal{X}-\mathcal{Y})}{\partial\alpha_{j}}=\mathrm{d}(\mathcal{X}-\mathcal{Y})=0\,.$ (12)
Picking one of the variables (say, without loss of generality, $\alpha_{0}$) appearing in $\mathcal{X}-\mathcal{Y}$, we can express the integration measure $\omega$ on $S_{n}$ as
$\left.\omega\right|_{S_{n}}$ $=\left|\frac{\partial\left(\mathcal{X}-\mathcal{Y}\right)}{\partial\alpha_{0}}\right|_{S_{n}}^{-1}\left.\mathcal{X}\right|_{S_{n}}\mathrm{d}\alpha_{1}\wedge\ldots\wedge\mathrm{d}\alpha_{n}$
$=\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\,\mathrm{d}\alpha_{0}\wedge\mathrm{d}\alpha_{1}\wedge\ldots\wedge\mathrm{d}\alpha_{n}\,.$ (13)
This then leads to Eq. (11).
Note that $f$ is a homogeneous function of degree $(-n-1)$ in the variables $\{\alpha_{j}\}$. This can be used to demonstrate that the $n$-form $f\,\omega$ is closed, i.e., $\mathrm{d}(f\,\omega)=\mathrm{d}f\wedge\omega+f\,\mathrm{d}\omega=0$:
$\mathrm{d}f\wedge\omega$ $=\left(\sum_{j=0}^{n}\frac{\partial f}{\partial\alpha_{j}}\mathrm{d}\alpha_{j}\right)\wedge\left(\sum_{j=0}^{n}(-1)^{j}\,\alpha_{j}\,\mathrm{d}\alpha_{0}\wedge\ldots\wedge\widehat{\mathrm{d}\alpha_{j}}\wedge\ldots\wedge\mathrm{d}\alpha_{n}\right)$
$=-(n+1)\,f\,\mathrm{d}\alpha_{0}\wedge\mathrm{d}\alpha_{1}\wedge\ldots\wedge\mathrm{d}\alpha_{n}$
$=-f\,\mathrm{d}\omega\,.$ (14)
Therefore, the integration of $\mathrm{d}(f\,\omega)$ in any region $\Delta_{n+1}$ in $\mathbb{R}^{n+1}$ is zero. By Stokes’ theorem, this means
$\iint_{\Delta_{n+1}}\mathrm{d}(f\,\omega)=\oint_{\partial\Delta_{n+1}}f\,\omega=0\,.$ (15)
Now, suppose that $S_{n}$ and $S_{n}^{\prime}$ are two $n$-dimensional hypersurfaces defined by $\mathcal{X}-\mathcal{Y}=0$ and $\mathcal{X}^{\prime}-\mathcal{Y}^{\prime}=0$, respectively, and let $\Delta_{n+1}$ be the region enclosed by $S_{n}$, $S_{n}^{\prime}$ and the coordinate hyperplanes defined by $\alpha_{j}=0$ ($j=0,1,\ldots,n$). When restricted on the coordinate hyperplanes, $\omega$ always vanishes due to the factor of $\alpha_{j}$ in its definition (9). It is possible that we need to add hyperplanes at infinity to form a closed region. However, the integral at infinity does not contribute in dimensional regularization. Therefore, we can employ Eq. (15) to show that the integral is independent of the integration domain:
$\int_{S_{n}}f\,\omega=\int_{S_{n}^{\prime}}f\,\omega\,.$ (16)
Figure 1 illustrates the above idea for the case of two variables.
Figure 1: The independence of the integral on the integration domain for the case of two variables. $S_{1}$ is the integration domain defined by $\mathcal{X}=1$ and $\mathcal{Y}=\alpha_{0}+\alpha_{1}$, while $S_{1}^{\prime}$ is the integration domain defined by two functions $\mathcal{X}^{\prime}(\alpha_{0},\alpha_{1})$ and $\mathcal{Y}^{\prime}(\alpha_{0},\alpha_{1})$.
In the above, we have shown that the two contours $S_{1}$ and $S^{\prime}_{1}$ give rise to the same integral, and can be regarded as belonging to the same equivalence class. In the language of twisted homology (see, e.g., Aomoto:1414035 for more details), the equivalence classes of integration contours are elements (cycles) of a twisted homology group determined by the polynomial $\alpha_{0}\,\mathcal{U}+\mathcal{F}$. This homology group is dual to the twisted cohomology group of the integrands. In the literature Mizera:2017rqa ; Mastrolia:2018uzb , there have been extensive discussions on how to perform integral reduction using the vector-space structure of the cohomology groups. This can be done using the techniques of intersection theory zbMATH00713739 ; zbMATH01270294 ; zbMATH02112802 ; zbMATH06267077 ; zbMATH06447521 ; zbMATH06502598 ; zbMATH06454357 ; zbMATH07531013 ; zbMATH07527773 ; zbMATH07733418 . The homology groups of integration contours also admit a vector-space structure, and in principle can be used for integral reduction as well. However, this path has not been followed in the literature to the best of our knowledge.
In the following, we will exploit the equivalence of integration domains to set up recursion relations that can be used to reduce one-loop integrals. We will write the relations in terms of index raising and lowering operators defined as:
$\hat{j}^{+}I(\cdots,\nu_{j},\cdots)$ $\equiv\nu_{j}I(\cdots,\nu_{j}+1,\cdots)\,,$
$\hat{j}^{-}I(\cdots,\nu_{j},\cdots)$ $\equiv I(\cdots,\nu_{j}-1,\cdots)\,.$ (17)
It is also useful to define an operator that set an index to zero:
$\hat{j}_{0}I(\cdots,\nu_{j},\cdots)\equiv I(\cdots,0,\cdots)\,.$ (18)
Before going into more general formalities, we first study a few simple examples.
2.2 A reducible sector
Figure 2: The massless triangle with $p_{1}^{2}=s$ and $p_{2}^{2}=p_{3}^{2}=0$.
To illustrate our approach, we first consider a simple case where a sector is reducible to one of its sub-sectors. A specific example is the massless triangle diagram shown in Figure 2. The corner integral in this sector corresponding to $\bm{\nu}_{3}=\bm{1}_{3}=\{1,1,1\}$ is given by
$I({\bm{1}}_{3})=C({\bm{1}}_{3})\int_{0}^{\infty}\mathrm{d}\alpha_{0}\mathrm{d}\alpha_{1}\mathrm{d}\alpha_{2}\mathrm{d}\alpha_{3}\,\alpha_{0}^{d-4}\left[\alpha_{0}\,\mathcal{U}+\mathcal{F}\right]^{-d/2}\mathcal{X}\,\delta(\mathcal{X}-\mathcal{Y})\,,$ (19)
where $\mathcal{U}=\alpha_{1}+\alpha_{2}+\alpha_{3}$ and $\mathcal{F}=-\alpha_{1}\alpha_{2}s$. The above integral is reducible to a bubble integral. One can observe that $\alpha_{3}$ is present in $\mathcal{U}$ but absent in $\mathcal{F}$. We will show that there is a general rule: whenever a variable is present in $\mathcal{U}$ but absent in $\mathcal{F}$, the corresponding integral is reducible to the sub-sector without the variable.
Let’s consider an $n$-point one-loop integral, and denote such a variable as $\alpha_{l}$. Defining $\mathcal{U}_{l,0}\equiv\mathcal{U}\big{|}_{\alpha_{l}\to 0}$, we can write $\mathcal{U}=\alpha_{l}+\mathcal{U}_{l,0}$. The corner integral in this sector can then be written as
$I({\bm{1}}_{n})=C({\bm{1}}_{n})\int_{0}^{\infty}\left(\prod_{j=0}^{n}\mathrm{d}\alpha_{j}\right)\alpha_{0}^{d-n-1}\left[\alpha_{0}\left(\alpha_{l}+\mathcal{U}_{l,0}\right)+\mathcal{F}\right]^{\lambda_{0}}\mathcal{X}\,\delta(\mathcal{X}-\mathcal{Y})\,.$ (20)
We choose
$\mathcal{X}=\frac{\mathcal{U}_{l,0}}{\alpha_{l}+\mathcal{U}_{l,0}}\,,\quad\mathcal{Y}=\alpha_{0}\,,$ (21)
and integrate out $\alpha_{0}$ using the $\delta$-function. After that, the variable $\alpha_{l}$ only appears in a power of $\alpha_{l}+\mathcal{U}_{l,0}$. The integration over $\alpha_{l}$ can then be performed using the formula
$\int_{0}^{\infty}\mathrm{d}x\,x^{\beta-1}{\left(1+\frac{x}{\Lambda}\right)}^{\gamma-1}\,\theta\left(1+\frac{x}{\Lambda}\right)=\begin{cases}\Lambda^{\beta}B(\beta,1-\beta-\gamma)\,,&(\Lambda>0)\,,\\ \left(-\Lambda\right)^{\beta}B(\beta,\gamma)\,,&(\Lambda<0)\,.\end{cases}$ (22)
where the Beta function is
$B(z_{1},z_{2})=\frac{\Gamma(z_{1})\,\Gamma(z_{2})}{\Gamma(z_{1}+z_{2})}\,.$ (23)
Note that the two expressions for $\Lambda>0$ and $\Lambda<0$ are actually equivalent if $\beta\in\mathbb{Z}$.
After integrating over $\alpha_{l}$, we arrive at
$I({\bm{1}}_{n})=C({\bm{1}}_{n})\int_{0}^{\infty}\left(\prod_{j\neq 0,l}\mathrm{d}\alpha_{j}\right)\mathcal{U}_{l,0}\left(\mathcal{U}_{l,0}+\mathcal{F}\right)^{\lambda_{0}}\,.$ (24)
As promised, the integral now manifestly has the form of the LP representation in the sub-sector without $\alpha_{l}$. Using the index-changing operators, the above result can be written as
$I({\bm{1}}_{n})=\frac{1}{d-n-1}\sum_{k\neq l}\hat{k}^{+}\hat{l}_{0}I(\bm{1}_{n})\,.$ (25)
Applying the above general formula to the triangle integral in Figure 2, we have
$I\left(1,1,1\right)=\frac{1}{d-4}\left[I\left(2,1,0\right)+I\left(1,2,0\right)\right].$ (26)
2.3 The bubble family
Figure 3: The bubble diagram with two masses.
We now turn to the reduction in an irreducible sector: the one-loop bubble with two masses shown in Figure 3. The kinematic variables are $m_{1}^{2}$, $m_{2}^{2}$ and $p^{2}=s$. The integrals in this family can be represented as
$I(\nu_{1},\nu_{2})=C(\nu_{1},\nu_{2})\int_{0}^{\infty}\mathrm{d}\alpha_{0}\mathrm{d}\alpha_{1}\mathrm{d}\alpha_{2}\,\alpha_{0}^{\nu_{0}-1}\alpha_{1}^{\nu_{1}-1}\alpha_{2}^{\nu_{2}-1}\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\\ \times\left[(\alpha_{1}+\alpha_{2})\alpha_{0}+\alpha_{1}\alpha_{2}(m_{1}^{2}+m_{2}^{2}-s)+\alpha_{1}^{2}m_{1}^{2}+\alpha_{2}^{2}m_{2}^{2}\right]^{\lambda_{0}}\,.$ (27)
We consider the reduction of $I(1,2)$. For that we introduce an auxiliary integral
$g(1,1)\equiv C(1,1)\int_{0}^{\infty}\mathrm{d}\alpha_{0}\mathrm{d}\alpha_{1}\mathrm{d}\alpha_{2}\ \alpha_{0}^{d-3}\mathcal{X}\delta\left(\mathcal{X}-\mathcal{Y}\right)\left\{\left[\alpha_{1}+(1-q_{1})\alpha_{2}\right]\alpha_{0}+\alpha_{1}^{2}m_{1}^{2}\right\}^{\lambda_{0}}\,,$ (28)
where $q_{1}\equiv(m_{1}^{2}+m_{2}^{2}-s)/(2m_{1}^{2})$. According to the result in Section 2.2, one can see that $g(1,1)$ is reducible to the sub-sector without $\alpha_{2}$. Applying the method from Section 2.2, we obtain
$g(1,1)=\frac{1}{d-3}\left(\frac{2m_{1}^{2}}{m_{1}^{2}-m_{2}^{2}+s}\right)I(2,0)\,.$ (29)
Alternatively, we can make a variable change $\alpha_{1}\to\alpha_{1}+q_{1}\alpha_{2}$ and $\alpha_{2}\to\alpha_{2}(1-W\alpha_{2}/\alpha_{0})$ in Eq. (28), where $W\equiv\lambda(m_{1}^{2},m_{2}^{2},s)/\left[2(m_{1}^{2}-m_{2}^{2}+s)\right]$ and the Källén function is given by $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2yz-2zx$. We can then arrive at
$g(1,1)$ $=C(1,1)\int_{0}^{\infty}\mathrm{d}\alpha_{0}\mathrm{d}\alpha_{2}\int_{-q_{1}\alpha_{2}}^{\infty}\mathrm{d}\alpha_{1}\,\alpha_{0}^{d-3}\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\left(1-2W\frac{\alpha_{2}}{\alpha_{0}}\right)$
×$\quad\times\left[(\alpha_{1}+\alpha_{2})\alpha_{0}+\alpha_{1}\alpha_{2}(m_{1}^{2}+m_{2}^{2}-s)+\alpha_{1}^{2}m_{1}^{2}+\alpha_{2}^{2}m_{2}^{2}\right]^{\lambda_{0}}$
$=I(1,1)+\frac{2W}{d-3}\,I(1,2)+h_{1}(1,1)\,,$ (30)
where we have split the integration range of $\alpha_{1}$ and define
$h_{1}(1,1)$ $\equiv C(1,1)\int_{0}^{\infty}\mathrm{d}\alpha_{0}\mathrm{d}\alpha_{2}\int_{-q_{1}\alpha_{2}}^{0}\mathrm{d}\alpha_{1}\,\alpha_{0}^{d-3}\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\left(1-2W\frac{\alpha_{2}}{\alpha_{0}}\right)$
×$\quad\times\left[(\alpha_{1}+\alpha_{2})\alpha_{0}+\alpha_{1}\alpha_{2}(m_{1}^{2}+m_{2}^{2}-s)+\alpha_{1}^{2}m_{1}^{2}+\alpha_{2}^{2}m_{2}^{2}\right]^{\lambda_{0}}\,.$ (31)
For the $h_{1}(1,1)$ function, we further apply the variable change $\alpha_{0}\to\alpha_{0}\alpha_{2}/(\alpha_{2}+\alpha_{1}/q_{1})$, $\alpha_{1}\to-\alpha_{1}\alpha_{2}/(\alpha_{2}+\alpha_{1}/q_{1})$, and denote $\alpha_{1}^{\prime}\equiv\alpha_{1}\left[1-W(2\alpha_{2}+\alpha_{1}/q_{1})/\alpha_{0}\right]$. We then obtain
$h_{1}(1,1)$ $=C(1,1)\int_{0}^{\infty}\mathrm{d}\alpha_{0}\mathrm{d}\alpha_{1}^{\prime}\mathrm{d}\alpha_{2}\,\alpha_{0}^{d-3}\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\left\{\left[(1-q_{1})\frac{\alpha_{1}^{\prime}}{q_{1}}+\alpha_{2}\right]\alpha_{0}+\alpha_{2}^{2}m_{2}^{2}\right\}^{\lambda_{0}}$
$=\frac{1}{d-3}\,\frac{m_{1}^{2}+m_{2}^{2}-s}{m_{1}^{2}-m_{2}^{2}+s}\,I(0,2)\,,$ (32)
where we have again utilized the method of Section 2.2. Combining Eq. (29), (2.3) and (2.3), we arrive at
$I(1,2)=(3-d)\,\frac{m_{1}^{2}-m_{2}^{2}+s}{\lambda(m_{1}^{2},m_{2}^{2},s)}\,I(1,1)+\frac{2m_{1}^{2}}{\lambda(m_{1}^{2},m_{2}^{2},s)}\,I(2,0)-\frac{m_{1}^{2}+m_{2}^{2}-s}{\lambda(m_{1}^{2},m_{2}^{2},s)}\,I(0,2)\,,$ (33)
which precisely agrees with the result of IBP reduction.
From the above examples, one may find that our procedure mainly consists of two steps: 1) transform of the integration contours either by explicit choices of the $\mathcal{X}$ and $\mathcal{Y}$ functions, or by appropriate variable changes; 2) split the integration contour into several parts, and identify each part manifestly as a Feynman integral. One may then wonder how these contour transforms are constructed. In the next Section, we will present the general method and the explicit recursive reduction formula for one-loop integrals.
3 The general method for one-loop integral reduction
3.1 More powers v.s. more variables
The goal of one-loop integral reduction is to reduce the indices $\nu_{j}$ to either 00 or $1$. We now introduce an interesting technique to lower an index by one, at the cost of adding an auxiliary integration variable. While this seems to be meaningless at first sight, it will be employed in the derivation of the final reduction formula.
For an arbitrary function $g(u)$, we can derive the following integral relation:
$\frac{1}{\Gamma(n+1)}\int_{0}^{\infty}\mathrm{d}u\,u^{n}g(u)$ $=\frac{1}{\Gamma(n+1)}\int_{0}^{\infty}\mathrm{d}u\left[n\int_{0}^{u}\mathrm{d}x\,x^{n-1}\right]g(u)$
$=\frac{1}{\Gamma(n)}\int_{0}^{\infty}\mathrm{d}x\int_{x}^{\infty}\mathrm{d}u\left[x^{n-1}g(u)\right]$
$=\frac{1}{\Gamma(n)\,\Gamma(1)}\int_{0}^{\infty}\mathrm{d}x\int_{0}^{\infty}\mathrm{d}y\left[x^{n-1}g(x+y)\right].$ (34)
Here, we emphasize that a regularization $n\to n+\rho$ is implicitly assumed if $n\leq 0$. We may apply the above relation to the Feynman parametric integrals. We define
$I(\nu_{1},\cdots,\left[\nu_{l},1\right],\cdots,\nu_{n})\\ \equiv C({\bm{\nu}}_{n},1)\int_{0}^{\infty}\mathrm{d}\beta_{l}\int_{0}^{\infty}\left(\prod_{j=0}^{n}\alpha_{j}^{\nu_{j}-1}\mathrm{d}\alpha_{j}\right)\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\left[\left({\alpha_{0}\,\mathcal{U}}+{\mathcal{F}}\right)\big{|}_{\alpha_{l}\to\alpha_{l}+\beta_{l}}\right]^{\lambda_{0}}\,.$ (35)
A few words are needed to explain a subtlety in the above definition. Both the prefactor $C$ and the power of $\alpha_{0}$ in the integrand involves $\nu$, which is defined as the sum of all indices in the argument of $I$ and $C$, as can be seen below Eq. (2). It then follows that, in the above expression, we have $\nu=1+\sum_{j=1}^{n}\nu_{j}$. Note that the value of $\nu_{0}\equiv-\nu-(L+1)\lambda_{0}$ is determined accordingly. Similar considerations apply for the functions $g(\bm{\nu})$ and $h_{i}(\bm{\nu})$ that will be introduced later.
Applying Eq. (3.1) to Eq. (35), we obtain
$I(\nu_{1},\cdots,\nu_{l}+1,\cdots,\nu_{n})=I(\nu_{1},\cdots,\left[\nu_{l},1\right],\cdots,\nu_{n})\,.$ (36)
From a different point of view, the above relation can also be obtained by splitting $D_{l}^{-\nu_{l}-1}$ as $D_{l}^{-\nu_{l}}\,D_{l}^{-1}$, and introducing two Feynman parameters $\alpha_{l}$ and $\beta_{l}$ for the two factors. This relation will play a crucial role in subsequent derivations. It is worth noting that it also holds for general $L$-loop integrals.
3.2 Reduction for irreducible sectors
We now consider a general one-loop integral in the Feynman representation:
$I({\bm{\nu}}_{n})=C({\bm{\nu}}_{n})\int_{0}^{\infty}\left(\prod_{j=0}^{n}\alpha_{j}^{\nu_{j}-1}\mathrm{d}\alpha_{j}\right)\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\left[\left(\bm{1}^{T}\bm{\alpha}\right)\alpha_{0}+\bm{\alpha}^{T}\bm{Z}\bm{\alpha}\right]^{\lambda_{0}}\,,$ (37)
where $\bm{1}$ denotes a column vector of length $n$ with all elements equal to $1$, and $\bm{\alpha}\equiv(\alpha_{1},\cdots,\alpha_{n})^{T}$. It is evident that
$\mathcal{U}=\bm{1}^{T}\bm{\alpha}\,,\quad\mathcal{F}=\bm{\alpha}^{T}\bm{Z}\bm{\alpha}\,,$ (38)
where the Gram matrix $\bm{Z}$ is symmetric. For an irreducible sector, we have $\det(\bm{Z})\neq 0$ (but the reverse is not true, as we will see in the “magic relations” discussed later).
In the following, we will aim for decreasing the power $\nu_{l}>1$ of the variable $\alpha_{l}$, where $l\in\{1,\cdots,n\}$. For that we will need to study the submatrix of $\bm{Z}$ obtained by removing the $l$-th row and $l$-th column. We will denote this submatrix as $\bm{Z}(\hat{l},\hat{l})$, and refer to it as “the $(l,l)$-submatrix”. A closely related concept is the $(i,j)$-minor of $\bm{Z}$, which is the determinant of the submatrix of $\bm{Z}$ without the $i$-th row and $j$-th column. The $(i,j)$-cofactor is further defined by multiplying the $(i,j)$-minor with $(-1)^{i+j}$. We will denote the $(i,j)$-cofactor of $\bm{Z}$ as $Z_{i,j}$.
The $(l,l)$-submatrix can be either non-singular with rank $n-1$, or singular with a smaller rank. In the following we will discuss the two situations separately.
3.2.1 When the $(l,l)$-submatrix is non-singular
Similar to the idea of Eq. (28), we introduce an auxiliary integral
$g({\bm{\nu}}_{n},{\bm{1}}_{n})\equiv C({\bm{\nu}}_{n},{\bm{1}}_{n})\int_{0}^{\infty}\left(\prod_{j=1}^{n}{\mathrm{d}\beta_{j}}\right)\int_{0}^{\infty}\left(\prod_{j=0}^{n}\alpha_{j}^{\nu_{j}-1}\,\mathrm{d}\alpha_{j}\right)\,\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\\ \times\left\{\left[\bm{1}^{T}\left(\bm{{\alpha}}+{\bm{\beta}}^{\left(l\right)}-\bm{q}\beta_{l}\right)\right]\alpha_{0}+{\left(\bm{\alpha}+{\bm{\beta}}^{\left(l\right)}\right)}^{T}\bm{Z}\left(\bm{\alpha}+{\bm{\beta}}^{\left(l\right)}\right)\right\}^{\lambda_{0}}\,,$ (39)
where $\bm{\beta}^{\left(l\right)}$ denotes the vector obtained by replacing the $l$-th element (i.e., $\beta_{l}$) of the vector $\bm{\beta}\equiv(\beta_{1},\cdots,\beta_{n})^{T}$ with 00, and the elements of the vector $\bm{q}$ are given by
$q_{k}\equiv-\frac{Z_{k,l}}{Z_{l,l}}\,,$ (40)
where we recall that $Z_{i,j}$ is the $(i,j)$-cofactor of $\bm{Z}$.
Using the method from Section 2.2, we can extract $\beta_{l}$ and integrate it out with an appropriate choice of $\mathcal{X}$ and $\mathcal{Y}$. We arrive at
$g({\bm{\nu}}_{n},{\bm{1}}_{n})=\frac{1}{\nu_{0}-1}\left(-\frac{1}{q}\right)\sum_{j=1}^{n}\left[\hat{j}^{+}\hat{l}^{-}I({\bm{\nu}}_{n}+{\bm{1}}_{n})\right]\,,$ (41)
where $q\equiv\sum_{k=1}^{n}q_{k}$, and we have used Eq. (36) to absorb the extra $\beta_{j}$ variables at the cost of increasing the powers of $\alpha_{j}$. Note that $\nu_{0}$ is defined in terms of the sequence of indices $(\bm{\nu}_{n},\bm{1}_{n})$, as explained below Eq. (35).
On the other hand, Eq. (39) can be transformed in another way with the variable change
$\beta_{j}\to\beta_{j}+q_{j}\beta_{l}\,,\;\left(\forall j\neq l\right),\quad\beta_{l}\to\beta_{l}\left(1-W\,\frac{\beta_{l}+2\alpha_{l}}{\alpha_{0}}\right),$ (42)
where
$W\equiv-\frac{\det(\bm{Z})}{\sum_{j=1}^{n}{Z_{j,l}}}=\frac{\det(\bm{Z})}{q\,Z_{l,l}}\,.$ (43)
This leads to
$g({\bm{\nu}}_{n},{\bm{1}}_{n})=C({\bm{\nu}}_{n},{\bm{1}}_{n})\int_{0}^{\infty}\mathrm{d}\beta_{l}\left(\prod_{j\neq l}\int_{-q_{j}\beta_{l}}^{\infty}{\mathrm{d}\beta_{j}}\right)\int_{0}^{\infty}\left(\prod_{j=0}^{n}\alpha_{j}^{\nu_{j}-1}\,\mathrm{d}\alpha_{j}\right)\,\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\\ \times\left(1-2W\frac{\alpha_{l}+\beta_{l}}{\alpha_{0}}\right)\left[\bm{1}^{T}\left(\bm{\alpha}+\bm{\beta}\right)\alpha_{0}+{\left(\bm{\alpha}+\bm{\beta}\right)}^{T}\bm{Z}{\left(\bm{\alpha}+\bm{\beta}\right)}\right]^{\lambda_{0}}\,.$ (44)
We can now split the integration domain of $\beta_{j}$ according to:
$\prod_{j\neq l}{\int_{-x_{l}}^{\infty}\mathrm{d}x_{j}}=\prod_{j\neq l}{\int_{0}^{\infty}\mathrm{d}x_{j}}+\sum_{k\neq l}\int_{-x_{l}}^{0}\mathrm{d}x_{k}\prod_{j\neq l,k}{\int_{x_{k}}^{\infty}\mathrm{d}x_{j}}\,.$ (45)
We then arrive at
$g({\bm{\nu}}_{n},{\bm{1}}_{n})=I({\bm{\nu}}_{n}+{\bm{1}}_{n})+\frac{2W}{\nu_{0}-1}\hat{l}^{+}I({\bm{\nu}}_{n}+{\bm{1}}_{n})+\sum_{k\neq l}h_{k}({\bm{\nu}}_{n},{\bm{1}}_{n})\,,$ (46)
where we have defined
$h_{k}({\bm{\nu}}_{n},{\bm{1}}_{n})\equiv C({\bm{\nu}}_{n},{\bm{1}}_{n})\int_{0}^{\infty}\mathrm{d}\beta_{l}\int_{-q_{k}\beta_{l}}^{0}\mathrm{d}\beta_{k}\left(\prod_{j\neq l,k}{\int_{q_{j}\beta_{k}/q_{k}}^{\infty}\mathrm{d}\beta_{j}}\right)\int_{0}^{\infty}\left(\prod_{j=0}^{n}\alpha_{j}^{\nu_{j}-1}\,\mathrm{d}\alpha_{j}\right)\\ \times\left(1-2W\frac{\alpha_{l}+\beta_{l}}{\alpha_{0}}\right)\left[\bm{1}^{T}\left(\bm{\alpha}+\bm{\beta}\right)\alpha_{0}+{\left(\bm{\alpha}+\bm{\beta}\right)}^{T}\bm{Z}{\left(\bm{\alpha}+\bm{\beta}\right)}\right]^{\lambda_{0}}\,\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\,.$ (47)
Our next task is to transform $h_{k}({\bm{\nu}}_{n})$ into Feynman integrals. For that we perform the variable change
$\alpha_{j}\to\alpha_{j}\frac{\beta_{l}}{\beta_{l}+\frac{\beta_{k}}{q_{k}}}\,,\;\left(\forall j\right),\quad\beta_{k}\to-\beta_{k}\frac{\beta_{l}}{\beta_{l}+\frac{\beta_{k}}{q_{k}}}\,,\quad\beta_{j}\to\left(\beta_{j}-q_{j}\frac{\beta_{k}}{q_{k}}\right)\frac{\beta_{l}}{\beta_{l}+\frac{\beta_{k}}{q_{k}}}\,,\ \left(\forall j\neq l,k\right).$ (48)
We can then write
$h_{k}({\bm{\nu}}_{n},{\bm{1}}_{n})=C({\bm{\nu}}_{n},{\bm{1}}_{n})\int_{0}^{\infty}\mathrm{d}\beta_{k}^{\prime}\int_{0}^{\infty}\left(\prod_{j\neq k}{\mathrm{d}\beta_{j}}\right)\int_{0}^{\infty}\left(\prod_{j=0}^{n}\alpha_{j}^{\nu_{j}-1}\,\mathrm{d}\alpha_{j}\right)\,\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\\ \quad\times\left\{\left[\bm{1}^{T}\left(\bm{\alpha}+{\bm{\beta}}^{\left(k\right)}\right)-q\frac{\beta_{k}^{\prime}}{q_{k}}\right]\alpha_{0}+\left(\bm{\alpha}+{\bm{\beta}}^{\left(k\right)}\right)^{T}\bm{Z}\left(\bm{\alpha}+{\bm{\beta}}^{\left(k\right)}\right)\right\}^{\lambda_{0}}\,,$ (49)
where
$\beta_{k}^{\prime}\equiv\beta_{k}\left(1-W\,\frac{2\alpha_{l}+2\beta_{l}+\beta_{k}/q_{k}}{\alpha_{0}}\right).$ (50)
Therefore
$h_{k}({\bm{\nu}}_{n},{\bm{1}}_{n})=\frac{1}{\nu_{0}-1}\left(-\frac{q_{k}}{q}\right)\sum_{j=1}^{n}\left[\hat{j}^{+}\hat{k}^{-}I({\bm{\nu}}_{n}+{\bm{1}}_{n})\right]\,.$ (51)
Combining Eq. (41), (46) and (51), we finally obtain
$2\det(\bm{Z})\,\hat{l}^{+}I(\bm{\nu}_{n})=\sum_{j=1}^{n}\left[{Z_{j,l}}(\nu_{0}-\nu_{j})\right]I(\bm{\nu}_{n})-\sum_{k=1}^{n}Z_{k,l}J_{k}({\bm{\nu}}_{n})\,.$ (52)
where we have performed a simple shift of indices $\nu_{i}\to\nu_{i}-1$, and
$J_{k}(\bm{\nu}_{n})\equiv\sum_{j\neq k}\hat{j}^{+}\hat{k}^{-}I(\bm{\nu}_{n})\,.$ (53)
In the context of integral reduction, we regard the integral which has a smaller total positive indices as “simpler”. In this sense, it can be seen that the integrals appearing on the right-hand side of Eq. (52) are simpler than the one on the left-hand side. Since we are dealing with an irreducible sector, we have $\det(\bm{Z})\neq 0$. Therefore, we can use Eq. (52) to recursively reduce a particular Feynman integral to simpler integrals.
3.2.2 When the $(l,l)$-submatrix is singular
In this case, the $(l,l)$-submatrix is non-invertible, which implies that $Z_{l,l}=0$. Eq. (40) then becomes singular. However, we find that the correct reduction rules can be obtained by taking the limit $Z_{l,l}\to 0$ in the final expression Eq. (52). Therefore, in this case, we have
$2\det(\bm{Z})\,\hat{l}^{+}I(\bm{\nu}_{n})=\sum_{j\neq l}\left[{Z_{j,l}}(\nu_{0}-\nu_{j})\right]I(\bm{\nu}_{n})-\sum_{k\neq l}\left[Z_{k,l}J_{k}({\bm{\nu}}_{n})\right],$ (54)
which can be used to reduce $\hat{l}^{+}I(\bm{\nu}_{n})$.
A special case occurs when the matrix $\bm{Z}$ has only one element, meaning that $\bm{Z}(\hat{l},\hat{l})$ is an empty matrix. In this case, we have $\nu=\nu_{1}$ and $\nu_{0}=d-\nu$. The procedure can stilled be carried out by defining $Z_{1,1}=1$, and we obtain the correct reduction formula for single-propagator integrals:
$I(\nu+1)=\frac{d-2\nu}{2\nu\,\det(\bm{Z})}I(\nu)\,.$ (55)
3.3 The reducible sectors
We now turn to the case where $\det(\bm{Z})=0\,$, which corresponds to reducible sectors. Noting that the Gram matrix $\bm{Z}$ is singular in this case, it implies that the degrees of freedom in the polynomial $\mathcal{F}$ are less than the number of variables. In another words, we can find a vector $\bm{\xi}=(\xi_{1},\cdots,\xi_{n})$ with $\xi_{l}=1$, such that
$\mathcal{F}(\alpha_{1}+\xi_{1}\alpha_{l},\ldots,\alpha_{l},\ldots,\alpha_{n}+\xi_{n}\alpha_{l})=\mathcal{F}(\alpha_{1},\ldots,0,\ldots,\alpha_{n})\,.$ (56)
The vector $\bm{\xi}$ is in the kernel of $\bm{Z}$, i.e., $\bm{Z}\bm{\xi}=\bm{0}$.
3.3.1 Reduction for general reducible sectors
In Section 2.2, we have discussed a special case where the $\mathcal{F}$ polynomial already takes the form of the right-hand side of Eq. (56), without introducing the transformation induced by $\bm{\xi}$. In other words, it corresponds to the case where all elements in $\bm{\xi}$, except $\xi_{l}=1$, are zero. In that case, integrals in this sector can be reduced to a single sub-sector. More generically, we need to apply the variable change
$\alpha_{j}\to\alpha_{j}+\xi_{j}\alpha_{l}\,,\quad(\forall j\neq 0,l)\,,$ (57)
to transform the $\mathcal{F}$ polynomial into the desired form. Consequently, the integrals will be reduced to linear combinations of different sub-sectors, as we will demonstrate in the following.
After the variable change, an integral can be written as
$I({\bm{\nu}}_{n})$ $=C({\bm{\nu}}_{n})\int_{0}^{\infty}\left(\prod_{j=0}^{n}\alpha_{j}^{\nu_{j}-1}\,\mathrm{d}\alpha_{j}\right)\,\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\left[\alpha_{0}\,\mathcal{U}+\mathcal{F}\right]^{\lambda_{0}}$
$=C({\bm{\nu}}_{n})\int_{0}^{\infty}\alpha_{l}^{\nu_{l}-1}\,\mathrm{d}\alpha_{l}\left(\prod_{j\neq 0,l}\int_{-\xi_{j}\alpha_{l}}^{\infty}\left(\alpha_{j}+\xi_{j}\alpha_{l}\right)^{\nu_{j}-1}\,\mathrm{d}\alpha_{j}\right)$
×$\quad\times\int_{0}^{\infty}\mathrm{d}\alpha_{0}\,\alpha_{0}^{\nu_{0}-1}\left[\left(\mathcal{U}_{l,0}+\xi\alpha_{l}\right)\alpha_{0}+\mathcal{F}_{l,0}\right]^{\lambda_{0}}\,\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\,,$ (58)
where $\xi\equiv\sum_{j=1}^{n}\xi_{j}$, $\mathcal{F}_{l,0}\equiv\mathcal{F}\big{|}_{\alpha_{l}\to 0}$ and we recall that $\mathcal{U}_{l,0}\equiv\mathcal{U}\big{|}_{\alpha_{l}\to 0}$. We can again split the integration domain using Eq. (45), and define
$G_{l}({\bm{\nu}}_{n})\equiv C({\bm{\nu}}_{n})\int_{0}^{\infty}\mathrm{d}\alpha_{l}\,\alpha_{l}^{\nu_{l}-1}\int_{0}^{\infty}\mathrm{d}\alpha_{0}\,\alpha_{0}^{\nu_{0}-1}\,\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\\ \times\int_{0}^{\infty}\left(\prod_{j\neq 0,l}\mathrm{d}\alpha_{j}\left(\alpha_{j}+\xi_{j}\frac{\alpha_{l}}{\xi_{l}}\right)^{\nu_{j}-1}\right)\left[\left(\mathcal{U}_{l,0}+\xi\frac{\alpha_{l}}{\xi_{l}}\right)\alpha_{0}+\mathcal{F}_{l,0}\right]^{\lambda_{0}}\,,$ (59)
and
$H_{k}({\bm{\nu}}_{n})\equiv C({\bm{\nu}}_{n})\int_{0}^{\infty}\mathrm{d}\alpha_{l}\,\alpha_{l}^{\nu_{l}-1}\int_{-\xi_{k}\alpha_{l}}^{0}{\mathrm{d}\alpha_{k}}\left(\alpha_{k}+\xi_{k}\alpha_{l}\right)^{\nu_{k}-1}\int_{0}^{\infty}\mathrm{d}\alpha_{0}\,\alpha_{0}^{\nu_{0}-1}\,\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\\ \times\left(\prod_{j\neq 0,l,k}\int_{\xi_{j}\alpha_{k}/\xi_{k}}^{\infty}{\mathrm{d}\alpha_{j}}\left(\alpha_{j}+\xi_{j}\alpha_{l}\right)^{\nu_{j}-1}\right)\left[\left(\mathcal{U}_{l,0}+\xi\alpha_{l}\right)\alpha_{0}+\mathcal{F}_{l,0}\right]^{\lambda_{0}}\,.$ (60)
We can then write
$I({\bm{\nu}}_{n})=G_{l}({\bm{\nu}}_{n})+\sum_{k\neq l}H_{k}({\bm{\nu}}_{n})\,.$ (61)
We can further show that $H_{k}({\bm{\nu}}_{n})$ can actually be rewritten as $G_{k}({\bm{\nu}}_{n})$. For that we introduce the variable change
$\alpha_{0}\to\alpha_{0}\,\frac{\alpha_{l}}{\alpha_{l}-\alpha_{k}/\xi_{k}}\,,\quad\alpha_{k}\to-\xi_{k}\alpha_{l}\,\frac{\alpha_{l}}{\alpha_{l}-\alpha_{k}/\xi_{k}}\,,$
$\alpha_{j}\to\left(\alpha_{j}-\xi_{j}\alpha_{l}\right)\frac{\alpha_{l}}{\alpha_{l}-\alpha_{k}/\xi_{k}}\,,\quad(\forall j\neq 0,k,l)\,,$ (62)
and therefore
$H_{k}({\bm{\nu}}_{n})=C({\bm{\nu}}_{n})\int_{0}^{\infty}\mathrm{d}\alpha_{k}\,\alpha_{k}^{\nu_{k}-1}\int_{0}^{\infty}\mathrm{d}\alpha_{0}\,\alpha_{0}^{\nu_{0}-1}\,\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\\ \times\int_{0}^{\infty}\left(\prod_{j\neq 0,k}\mathrm{d}\alpha_{j}\,\left(\alpha_{j}+\xi_{j}\frac{\alpha_{k}}{\xi_{k}}\right)^{\nu_{j}-1}\right)\left[\left(\mathcal{U}_{k,0}+\xi\frac{\alpha_{k}}{\xi_{k}}\right)\alpha_{0}+\mathcal{F}_{k,0}\right]^{\lambda_{0}}\,.$ (63)
Using the method in Section 2.2, we can reduce $G_{k}({\bm{\nu}}_{n})$ (for both $k=l$ and $k\neq l$) to integrals in a sub-sector. Combining these results, we arrive at the reduction of $I({\bm{\nu}}_{n})$:
$I({\bm{\nu}}_{n})=\sum_{k\in S_{\xi}}\frac{\left(\xi_{k}\right)^{\nu_{k}}}{\left(\nu_{k}-1\right)!}\hat{k}_{0}\hat{P}_{k}I({\bm{\nu}}_{n})\,,$ (64)
where $S_{\xi}$ is the subset of $\{1,\cdots,n\}$ such that $\xi_{k}\neq 0$ if $k\in S_{\xi}$, and the operator $\hat{P}_{k}$ is a combination of index-raising and lowering operators:
$\hat{P}_{k}=\left(\prod_{j\in S_{\xi}}^{j\neq k}\sum_{\mu_{j}=0}^{\nu_{j}-1}\right)\left[B(-\nu_{0}+1,\mu+\nu_{k})\left(\frac{\sum_{i\neq k}\hat{i}^{+}}{-\xi}\right)^{\mu+\nu_{k}}\prod_{m\in S_{\xi}}^{m\neq k}\frac{\left(\xi_{m}\hat{m}^{-}\right)^{\mu_{t}}}{\mu_{t}!}\right]\,,$ (65)
where $\mu\equiv\sum_{j\in S_{\xi}}^{j\neq k}\mu_{j}$. Obviously, the above formula only applies if $\xi\neq 0$. The case for $\xi=0$ will be discussed in the following.
3.3.2 Degenerate limits
The case $\xi=0$ corresponds to certain degenerate limits of Feynman integrals. As a specific example, consider the bubble diagram shown in Figure 3 with $m_{1}^{2}=m_{2}^{2}$ and $s=0$. Here we find that $\bm{\xi}=\{1,-1\}$, and hence $\xi=0$. We now discuss how to deal with this kind of situations.
When $\xi=0\,$, both $\mathcal{U}$ and $\mathcal{F}$ satisfy
$\mathcal{O}(\alpha_{1},\cdots,\alpha_{l},\cdots,\alpha_{n})=\mathcal{O}(\alpha_{1}-\xi_{1}\alpha_{l},\cdots,0,\cdots,\alpha_{n}-\xi_{n}\alpha_{l})\,,$ (66)
for certain $l$, where $\mathcal{O}\in\{\mathcal{U},\mathcal{F}\}$. Therefore $\alpha_{0}\,\mathcal{U}+\mathcal{F}$ satisfies the above property as well. We now consider the integral $I(\bm{\nu}_{n}+\bm{1}_{n})$. By introducing the $\beta$-variables as in Section 3.1, we can write it as
$I({\bm{\nu}}_{n},{\bm{1}}_{n})=C({\bm{\nu}}_{n},{\bm{1}}_{n})\int_{0}^{\infty}\left(\prod_{j=1}^{n}{\mathrm{d}\beta_{j}}\right)\int_{0}^{\infty}\left(\prod_{j=0}^{n}\alpha_{j}^{\nu_{j}-1}\,\mathrm{d}\alpha_{j}\right){\mathcal{G}_{0}(\bm{\alpha},\bm{\beta})}^{\lambda_{0}}\,\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\,,$ (67)
where
$\mathcal{G}_{0}(\bm{\alpha},\bm{\beta})\equiv\left[\alpha_{0}\,\mathcal{U}+\mathcal{F}\right]_{\bm{\alpha}\to\bm{\alpha}+\bm{\beta}}=\left[\bm{1}^{T}\left(\bm{\alpha}+\bm{\beta}\right)\right]\alpha_{0}+{\left(\bm{\alpha}+\bm{\beta}\right)}^{T}\bm{Z}\left(\bm{\alpha}+\bm{\beta}\right).$ (68)
It follows from the property of $\mathcal{U}$ and $\mathcal{F}$ that
$\mathcal{G}_{0}(\bm{\alpha},\beta_{1},\cdots,\beta_{l},\cdots,\beta_{n})=\mathcal{G}_{0}({\bm{\alpha}},\beta_{1}-\xi_{1}\beta_{l},\cdots,0,\cdots,\beta_{n}-\xi_{n}\beta_{l})\,.$ (69)
We now show that this property can be used for reducing the integral to sub-sectors.
For convenience, we consider a function $g(y,\bm{x}_{n-1})$ which satisfies
$g(y,\bm{x}_{n-1})=g(0,\bm{x}_{n-1}+y\bm{\eta}_{n-1})\,,$ (70)
where $\bm{x}_{n-1}$ denotes the sequence of $n-1$ variables, and $\bm{\eta}_{n-1}$ is the sequence of $n-1$ constants. Using the relation
$\int_{0}^{\infty}\mathrm{d}x_{l}\int_{x_{l}}^{\infty}\left(\prod_{j\neq l}\mathrm{d}x_{j}\right)=\sum_{k\neq l}\int_{0}^{\infty}\mathrm{d}x_{k}\int_{x_{k}}^{\infty}\left(\prod_{j\neq k,l}\mathrm{d}x_{j}\right)\int_{0}^{x_{k}}\mathrm{d}x_{l}\,,$ (71)
we can show that
$\int_{0}^{\infty}\mathrm{d}y\int_{0}^{\infty}\left(\prod_{j=1}^{n-1}\mathrm{d}x_{j}\right)g(y,\bm{x}_{n-1})$ $=\int_{0}^{\infty}\mathrm{d}y\left(\prod_{j=1}^{n-1}\int_{y\eta_{j}}^{\infty}\mathrm{d}x_{j}^{\prime}\right)g(0,\bm{x}_{n-1}^{\prime})$
$=\sum_{k=1}^{n-1}\int_{0}^{\infty}\mathrm{d}x_{k}^{\prime}\left(\prod_{j=1}^{n-1}\int_{\eta_{j}x_{k}^{\prime}/\eta_{k}}^{\infty}\mathrm{d}x_{j}^{\prime}\right)\int_{0}^{x_{k}^{\prime}}\mathrm{d}y\,g(0,\bm{x}_{n-1}^{\prime})$
$=\sum_{k=1}^{n-1}\eta_{k}\int_{0}^{\infty}\mathrm{d}Y\int_{0}^{\infty}\left(\prod_{j\neq k}\mathrm{d}X_{j}\right)Y\,g(Y,\bm{X}_{n-1})\,,$ (72)
where $\bm{x}_{n-1}^{\prime}\equiv\bm{x}_{n-1}+y\bm{\eta}_{n-1}$, $Y\equiv x_{k}^{\prime}/\eta_{k}=x_{k}/\eta_{k}+y$, and $X_{j}\equiv x_{j}^{\prime}-\eta_{j}x_{k}^{\prime}/\eta_{k}=x_{j}-\eta_{j}x_{k}/\eta_{k}$. Note that in the $k$th term of the above result, we have $X_{k}=0$ in $g$, and we don’t have to integrate over $X_{k}$. Hence, we have eliminated one variable at the cost of increasing the power of $Y$.
Applying Eq. (3.3.2) to $I({\bm{\nu}}_{n},{\bm{1}}_{n})$, we get
$I({\bm{\nu}}_{n},{\bm{1}}_{n})=-\sum_{k\neq l}\xi_{k}C({\bm{\nu}}_{n},{\bm{1}}_{n})\int_{0}^{\infty}\left(\prod_{j\neq k}{\mathrm{d}\beta_{j}}\right)\int_{0}^{\infty}\left(\prod_{j=0}^{n}\alpha_{j}^{\nu_{j}-1}\,\mathrm{d}\alpha_{j}\right)\\ \times\beta_{l}\,{\mathcal{G}_{0}(\bm{\alpha},{\bm{\beta}}^{\left(k\right)})}^{\lambda_{0}}\,\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\,.$ (73)
Absorbing the extra $\beta_{j}$ variables using Eq. (36), we finally obtain the reduction formula that works in the degenerate limits:
$I(\cdots,\nu_{l},\cdots)=-\sum_{k\neq l}\xi_{k}I(\cdots,\nu_{k}-1,\cdots,\nu_{l}+1,\cdots)\,.$ (74)
It can be observed that by fixing $l$, we can recursively decrease other indices at the cost of increasing $\nu_{l}$. This eventually leads to integrals in the sub-sectors.
Now, let us consider the example mentioned earlier: the bubble diagram with $m_{1}^{2}=m_{2}^{2}$ and $s=0$. By choosing $l=1$ and repeatedly applying Eq. (74), we can arrive at
$I(\nu_{1},\nu_{2})=I(\nu_{1}+\nu_{2},0)\,.$ (75)
Up to now, we have exhausted all possibilities and have presented iterative formulas to reduce any one-loop integral with indices larger than one to linear combinations of simpler integrals. The remaining task is to reduce the negative indices corresponding to numerators in the momentum representation Eq. (1).
3.4 Reduction of integrals with numerators
If any index $\nu_{j}<0$, the corresponding $D_{j}$ is in the numerator of Eq. (1). Such integrals can always be reduced to integrals with $\nu_{j}=0$. In the following, we show how to achieve this reduction within our approach.
3.4.1 Rewriting the integrals in shifted dimensions
It is well-known that integrals with $\nu_{j}<0$ can be rewritten as integrals with $\nu_{j}=0$ in shifted spacetime dimensions. This can be easily seen in the LP representation:
$I^{(d)}({\bm{\nu}}_{n})=C^{(d)}({\bm{\nu}}_{n})\int_{0}^{\infty}\mathcal{G}^{-d/2}\prod_{j=1}^{n}\alpha_{j}^{\nu_{j}-1}{\mathrm{d}\alpha_{j}}\,,$ (76)
where $\mathcal{G}\equiv\mathcal{U}+\mathcal{F}$. We have used the superscript $(d)$ to specify the dependence on the spacetime dimension.
If $\nu_{j}<0$, we can increase $\nu_{j}$ with a simple integration-by-parts:
$\int_{0}^{\infty}\mathrm{d}\alpha_{j}\frac{\alpha_{j}^{\rho+\nu_{j}-1}}{\Gamma(\rho+\nu_{j})}\mathcal{G}^{-d/2}=\frac{d}{2}\int_{0}^{\infty}\mathrm{d}\alpha_{j}\frac{\alpha_{r}^{\rho+\nu_{j}}}{\Gamma(\rho+\nu_{j}+1)}\left(\frac{\partial\mathcal{G}}{\partial\alpha_{j}}\right)\mathcal{G}^{-(d+2)/2}\,,$ (77)
where we have explicitly introduced the regulator $\rho$ as discussed in Section 2.1. One may see that the right-hand side of the above equation corresponds to integrals in $d+2$ dimensions, but with an increased $\nu_{j}$ index. Note that $\mathcal{U}$ and $\mathcal{F}$ are homogeneous polynomials of degree $L$ and $L+1$, respectively. The above relation can be rewritten in terms of index-raising operators as
$I^{(d)}(\cdots,\nu_{j},\cdots)=(-1)^{L}\left[\left((L+1)\frac{d}{2}-\sum_{i=1}^{n}\nu_{i}\right)\,\hat{\mathcal{U}}_{j}^{+}-\hat{\mathcal{F}}_{j}^{+}\right]I^{(d+2)}(\cdots,\nu_{j}+1,\cdots)\,,$ (78)
where $\hat{\mathcal{O}}_{j}^{+}$ (for $\mathcal{O}\in\{\mathcal{U},\mathcal{F}\}$) represents an operator obtained by substituting each $\alpha_{k}$ in $\left(\partial\mathcal{O}/\partial\alpha_{j}\right)$ by $\hat{k}^{+}$. Note that Eq. (78) holds for general $L$-loop integrals, and for the purpose of this work, we can set $L=1\,$.
With Eq. (78), we can recursively increase a negative index to zero, at the cost of employing integrals in shifted dimensions. Let us consider a simple example: $I(-1,2)$ from the bubble integral family in Figure 3. Applying Eq. (78) with $j=1$, we obtain
$I^{(d)}(-1,2)=\left(1-d\right)I^{(d+2)}(0,2)+2\left(m_{1}^{2}+m_{2}^{2}-s\right)I^{(d+2)}(0,3)\,.$ (79)
3.4.2 Dimensional recurrence relations
We now need to transform the integrals in shifted dimensions to the $d$-dimensional ones. This is the so-called dimensional recurrence relations Tarasov:1996br ; Tarasov:1997kx ; Lee:2009dh ; Lee:2010ug . We show how such relations can be naturally derived within our approach.
In the Feynman parameterization, it is easy to increase the dimension:
$I^{(d)}({\bm{\nu}}_{n})=-\sum_{k=1}^{n}\hat{k}^{+}I^{(d+2)}({\bm{\nu}}_{n})\,.$ (80)
We need to find the inverse relation, to express $I^{(d+2)}$ in terms of $I^{(d)}$. Note that the indices here are all non-negative. Since we can apply the reduction rules that we have constructed in both $d$ and $d+2$ dimensions, we only need to find the relations among corner integrals. Combining the above equation with Eq. (52), we obtain
$2\det(\bm{Z})I^{(d)}(\bm{1}_{n})$ $=-2\det(\bm{Z})\sum_{l=1}^{n}\hat{l}^{+}I^{(d+2)}(\bm{1}_{n})$
$=\zeta\left(n-1-d\right)I^{(d+2)}(\bm{1}_{n})-\sum_{k=1}^{n}\zeta_{k}\hat{k}_{0}I^{(d)}(\bm{1}_{n})\,,$ (81)
where $\zeta_{k}\equiv\sum_{j=1}^{n}Z_{k,j}$, $\zeta\equiv\sum_{k=1}^{n}\zeta_{k}$, and we have used
$\sum_{j\neq k}\hat{j}^{+}\hat{k}_{0}I^{(d+2)}(\bm{1}_{n})=-\hat{k}_{0}I^{(d)}(\bm{1}_{n})\,.$ (82)
If $\zeta\neq 0$, Eq. (3.4.2) gives rise to the dimension recurrence relation that we will need:
$I^{(d+2)}(\bm{1}_{n})=\frac{2\det(\bm{Z})}{\zeta\left(n-1-d\right)}I^{(d)}(\bm{1}_{n})+\frac{1}{\zeta\left(n-1-d\right)}\sum_{k=1}^{n}\zeta_{k}\hat{k}_{0}I^{(d)}(\bm{1}_{n})\,.$ (83)
Combining with Eq. (78), we complete the reduction of integrals with numerators.
Applying the above relation, we can continue the reduction of $I(-1,2)$ starting from Eq. (79). We first reduce the right-hand side of Eq. (79) to corner integrals in $d+2$ dimensions, and obtain
$I^{(d)}(-1,2)=-\frac{d}{2m_{2}^{2}}\frac{2m_{2}^{2}-\left(d-2\right)\left(m_{1}^{2}-m_{2}^{2}-s\right)}{2m_{2}^{2}}I^{(d+2)}(0,1)\,.$ (84)
We then use Eq. (83) to derive (recall that we have defined $Z_{1,1}=1$ when the matrix $\bm{Z}$ has only one element)
$I^{(d+2)}(0,1)=-\frac{2m_{2}^{2}}{d}I^{(d)}(0,1)\,.$ (85)
Finally, we have
$I^{(d)}(-1,2)=\frac{2m_{2}^{2}-\left(d-2\right)\left(m_{1}^{2}-m_{2}^{2}-s\right)}{2m_{2}^{2}}I^{(d)}(0,1)\,.$ (86)
Note that if $\zeta=0$, Eq. (3.4.2) does not lead to a dimension recurrence, but gives rise to an interesting reduction rule of the corner integral to sub-sectors:
$I^{(d)}(\bm{1}_{n})=-\frac{1}{2\det(\bm{Z})}\sum_{k=1}^{n}\zeta_{k}\hat{k}_{0}I^{(d)}(\bm{1}_{n})\,.$ (87)
This is the so-called “magic relation”. In the usual IBP reduction, it can only be observed when IBP relations involving certain super-sectors are included Klappert:2020nbg ; Frellesvig:2018ymi . For example, for the bubble diagram in Figure 3, we find $\zeta=s$. Therefore, when $s=0$, the integral $I(1,1)$ can be reduced to sub-sectors:
$I^{(d)}(1,1)=\frac{1}{m_{1}^{2}-m_{2}^{2}}\left[I^{(d)}(1,0)-I^{(d)}(0,1)\right]\,.$ (88)
In Kira, this reduction rule can only be found when embedding the bubble integrals in a triangle super-sector.
4 Examples
The recursive formulas derived in the previous Section allow a straightforward computer implementation. We have written a proof-of-concept Mathematica code to automatically reduce one-loop integrals. We have tested the program with various examples and confirmed its correctness.
Figure 4: The pentagon diagram with three masses.
As a more complicated example, consider the pentagon integral family with three masses shown in Figure 4. We set all external momenta to be light-like. The kinematic variables are $m_{1}^{2}$, $m_{2}^{2}$, $m_{3}^{2}$, $(p_{1}+p_{2})^{2}=s_{12}$, $(p_{2}+p_{3})^{2}=s_{23}$, $(p_{3}+p_{4})^{2}=s_{34}$, $(p_{4}+p_{5})^{2}=s_{45}$ and $(p_{5}+p_{1})^{2}=s_{51}$. The matrix $\bm{Z}$ in this family is
$\begin{pmatrix}m_{1}^{2}&\frac{1}{2}\left(m_{1}^{2}+m_{2}^{2}\right)&\frac{1}{2}\left(m_{1}^{2}-s_{12}\right)&\frac{1}{2}\left(m_{1}^{2}+m_{3}^{2}-s_{45}\right)&\frac{1}{2}m_{1}^{2}\\ \frac{1}{2}\left(m_{1}^{2}+m_{2}^{2}\right)&m_{2}^{2}&\frac{1}{2}m_{2}^{2}&\frac{1}{2}\left(m_{2}^{2}+m_{3}^{2}-s_{23}\right)&\frac{1}{2}\left(m_{2}^{2}-s_{51}\right)\\ \frac{1}{2}\left(m_{1}^{2}-s_{12}\right)&\frac{1}{2}m_{2}^{2}&0&\frac{1}{2}m_{3}^{2}&-\frac{1}{2}s_{34}\\ \frac{1}{2}\left(m_{1}^{2}+m_{3}^{2}-s_{45}\right)&\frac{1}{2}\left(m_{2}^{2}+m_{3}^{2}-s_{23}\right)&\frac{1}{2}m_{3}^{2}&m_{3}^{2}&\frac{1}{2}m_{3}^{2}\\ \frac{1}{2}m_{1}^{2}&\frac{1}{2}\left(m_{2}^{2}-s_{51}\right)&-\frac{1}{2}s_{34}&\frac{1}{2}m_{3}^{2}&0\end{pmatrix}.$ (89)
We consider the reduction of $I(1,2,1,1,1)$. Since $\bm{Z}$ is non-singular, we should apply Eq. (52) with $\bm{\nu}_{n}=(1,1,1,1,1)$ and $l=2$, which leads to
[…]
where $\bm{M}$ represents the matrix formed by replacing all elements in the second row of $\bm{Z}$ with $1$. The sub-sector integrals in the above expression may still contain integrals that need to be reduced (e.g., $I(2,1,1,1,0)$, $I(0,1,2,1,1)$, etc.). We can apply the relevant reduction formulas recursively until only master integrals remain. We have verified that the results agree with those provided by Kira.
We can also consider a reducible case, with $m_{1}^{2}=m_{2}^{2}=m_{3}^{2}=s_{12}=0$. At this kinematic point, the $\bm{Z}$ matrix is singular. Therefore, we need to apply Eq. (64) when performing the reduction of $I(1,2,1,1,1)$. The $\bm{\xi}$ vector can be derived as
$\bm{\xi}=\left\{-\frac{s_{23}}{s_{45}},1,-\frac{s_{51}}{s_{34}},0,0\right\}.$ (91)
It follows that $I(1,2,1,1,1)$ can be reduced to integrals in 3 sub-sectors:
$I(1,2,1,1,1)$ $=\frac{1}{(d-7)(d-8)}\frac{2s_{34}^{2}s_{45}^{2}}{(s_{23}s_{34}-s_{34}s_{45}+s_{45}s_{51})^{2}}\,\big{[}I(2,0,2,1,1)+I(2,0,1,2,1)$
$\quad+I(2,0,1,1,2)+I(1,0,2,2,1)+I(1,0,2,1,2)+I(1,0,1,2,2)$
$\quad+I(3,0,1,1,1)+I(1,0,3,1,1)+I(1,0,1,3,1)+I(1,0,1,1,3)\big{]}$
[…]
where the other two sub-sectors are those where the first or the third propagator is removed (e.g., $I(0,1,1,2,2)$, $I(1,1,0,2,2)$, etc.).
Finally, we consider a degenerate limit, where $p_{1}=p_{2}$ in addition to $m_{1}^{2}=m_{2}^{2}=m_{3}^{2}=0$. In this case, we have $s_{12}=0$, $s_{23}=s_{45}/2$ and $s_{51}=s_{34}/2$. Clearly, the top-sector is reducible, but we cannot apply Eq. (64) because $\bm{\xi}=\{1,-2,1,0,0\}$, which implies $\xi=0$. Therefore, in this case, we need to apply Eq. (74). By choosing $l=1$, we ultimately obtain
$I(1,2,1,1,1)=4I(3,0,1,1,1)-2I(3,1,0,1,1)-I(2,2,0,1,1)\,.$ (93)
5 Summary and outlook
In this paper, we have explored interesting properties of the Feynman parameterization that are useful for Feynman integral reduction without employing IBP relations. In particular, we have derived an extension of the Cheng-Wu theorem, such that the delta-function in the Feynman parameterization can take a more general form. This can be regarded as a specific implementation of the equivalence of integration contours, which is related to the twisted homology of Feynman integrals. Leveraging these properties, we have derived universal recursive formulas for the reduction of one-loop integrals in both irreducible and reducible sectors, and our method works in degenerate limits as well. The reduction relations can be easily implemented in a computer program. We have applied them to various examples and observed remarkable performance.
Figure 5: The sunrise diagram with $p^{2}=s$.
Given the success in one-loop problems, it is natural to extend our approach to multi-loop integrals. A new ingredient at higher loops is that there can be more than one MIs within a sector, and it is no longer guaranteed that all indices can be reduced to $1$ or 00. Correspondingly, in our approach, we find that when we perform the variable transformations, certain variables appear in the denominators of the transformed integrands. Such transformations are therefore only valid if the corresponding indices are greater than $1$. We have made initial attempts in the sunrise families shown in Figure 5. In the massless case $m_{1}=m_{2}=m_{3}=0$, we can derive reduction relations such as
$I(1,2,2)=\frac{10-3d}{2s}\left[I(1,2,1)+I(1,1,2)\right].$ (94)
Note that in the massless case, $I(1,2,1)$ and $I(1,1,2)$ can be further reduced to $I(1,1,1)$. This is one of the degenerate limits, which have a much richer structure than the one-loop cases. It deserves future investigation how to handle all possible degenerate limits beyond one-loop.
An additional higher-loop feature is that the relations that we can derive by naively extending the one-loop method often involve more than one integrals with the highest complexity. This means that we cannot derive a reduction rule from a single equation, but need to combine several equations. As an example, for the all-massive sunrise family where $m_{1}$, $m_{2}$ and $m_{3}$ are different and larger than zero, we may derive equations such as
$2m_{1}^{2}\left[I(2,2,1)+I(2,1,2)\right]+2m_{2}^{2}\,I(1,3,1)+2m_{3}^{2}\,I(1,1,3)\\ +\left(m_{1}^{2}+m_{2}^{2}+m_{3}^{2}-s\right)I(1,2,2)=\frac{3d-10}{2}\left[I(1,2,1)+I(1,1,2)\right]-I(0,2,2)\,.$ (95)
By solving such a system of equations, we can obtain the desired reduction rules. Note that the number of equations generated in our approach is generically much smaller than that of the usual IBP relations. Furthermore, in our initial attempts we have only naively applied variable transformations similar to the one-loop method. There are still significant flexibility in the integration contours yet to be fully explored. We believe that the system of equations can be further simplified, and it remains possible that one may derive straightforward reduction rules for higher-loop integrals as well.
Finally, in this work we have only exploited the equivalence of integration contours in a simple way. There can be deeper mathematical structures behind these equivalence relations. For example, it is possible to employ the concept of intersection numbers between two integration contours to directly compute the reduction coefficients. This provides a particularly intriguing future perspective.
Acknowledgements.
The authors would like to thank Liangliang Zhao, Tingfei Li, Mingming Lu and Yiyang Zhang for useful discussions. This work was supported in part by the National Natural Science Foundation of China under Grant No. 12375097, 12347103, and the Fundamental Research Funds for the Central Universities.